With Ha : μ ≠ 83.1
you obtain a test statistic of z = 2.33
Find the p-value accurate to 3 decimal places.
Answer #1

solution

two tail test

Test statistic = z=2.33

P(z >2.33 ) = 1 - P(z <2.33 ) = 1-0.9901=0.0099 z table

P-value = 2*0.0099=0.0198

p value =0.020 rounded